OpenStudy (anonymous):

how would i write y-3=4(x+1) in standard form

OpenStudy (anonymous):

writing equation in standard form? y+2= -3/4 (x+1) Standard form consists of these 7 parts in this order: 1. A positive whole number (if possible) 2. The letter x multiplied by the number in 1 3. A plus or minus sign 4. A whole number (if possible) 5. The letter y multiplied by the number in 4 6. An equal sign 7. A whole number (if possible) y + 2 = -3/4(x + 1) Multiply both sides by 4 to clear of fractions: 4(y + 2) = 4[-3/4(x + 1]] Distibute the 4 on the left. Cancel the 4's on the right: 1 4y + 8 = 4[-3/4(x + 1]] 1 4y + 8 = -3(x + 1) Distribute the 3 on the right 4y + 8 = -3x - 3 Add 3x to both sides: 4y + 8 + 3x = -3x - 3 + 3x 4y + 8 + 3x = -3 Subtract 8 from both sides 4y + 8 + 3x - 8 = -3 - 8 4y + 3x = -11 Reverse the terms on the left: 3x + 4y = -11

OpenStudy (anonymous):

@sarimari thank you

OpenStudy (anonymous):

your welcome any other question ?

OpenStudy (anonymous):

not right now

OpenStudy (anonymous):

okay